Vegas Company has the following unit costs:
  

Variable manufacturing overhead $ 25
Direct materials 20
Direct labor 19
Fixed manufacturing overhead 12
Variable marketing and administrative 7


Vegas produced and sold 10,000 units. If the product sells for $100, what is the contribution margin?

A. $360,000

B. $240,000

C. $290,000

D. $170,000
Answer #1

Total variable cost per unit=Direct materials+Direct labor+Variable manufacturing overhead+Variable marketing and administrative

=(20+19+25+7)=$71 per unit

Hence contribution margin=Sale-Variable cost

=(100-71)=$29 per unit

=($29*10,000)

=$290,000.